# arXiv:2603.00918v1[cs.CV] 01 Mar 2026

### v1

@@ Abstract
Text-to-image generation powers content creation across design, media, and data augmentation. Post-training of text-to-image generative models is a promising path to better match human preferences, factuality, and improved aesthetics. We introduce ARC (Adaptive Rewarding by Self-Confidence), a post-training framework that replaces external reward supervision with an internal self-confidence signal, obtained by evaluating how accurately the model recovers injected noise under self-denoising probes. ARC converts this intrinsic signal into scalar rewards, enabling fully unsupervised optimization without additional datasets, annotators, or reward models. Empirically, by reinforcing high-confidence generations, ARC delivers consistent gains in compositional generation, text rendering and text-image alignment over the baseline. We also find that integrating ARC with external rewards results in a complementary improvement, with alleviated reward hacking.
Figure 1: Qualitative examples of ARC on Pick-a-Pic dataset [30]. Best viewed on electronics.
@@ 1 Introduction
Despite extensive progress in extrinsically supervised post-training, relying on intrinsic signals remains under-explored for text-to-image generation. In this work, we aim to answer a fundamental question: can internal feedback from the text-to-image generator itself provide meaningful signals for post-training? To this end, we introduce Adaptive Rewarding by Self-Confidence (ARC), a post-training framework that relies on intrinsic self-certainty as the model-native reward. Inspired by Score Distillation Sampling [50, 62, 74], which leverages a pretrained text-to-image generator as a critic for text-to-3D or -4D generation, we propose to leverage a text-to-image generator to critique its own generation. Concretely, given a sampled latent $z_{0}$, we re-noise $z_{0}$ to selected timesteps $t\in\mathcal{T}$ using the model’s forward noising schedule, and measure how well the model reconstructs the injected noise, rewarding small residuals as ’high confidence’. Our hypothesis is that large-scale pretraining in diffusion generative models endows them with priors over real images and text–image alignment, so the model’s self-confidence would align strongly with text-alignment and realism.
Empircally, ARC yields consistent gains in compositional generation [21], text rendering [14], and text–image alignment [53], while modestly improving human-preference scores [30, 75, 78, 72] without external rewards. Qualitative side-by-side comparisons and a comprehensive user study corroborate these trends. Together, the results indicate that intrinsic self-confidence aligns with key facets of image generation. Moreover, applying ARC on top of an extrinsically post-trained model (i.e., using external rewards) recovers additional improvements in compositionality, text rendering, and alignment, with only slight drops on the targeted external metric, evidencing that intrinsic and extrinsic rewards are complementarily beneficial for post-training stronger T2I generators.
@@
  • We present ARC (Adaptive Rewarding by self-Confidence), a post-training framework to use self-confidence as a reward for text-to-image generators.
  • We define ARC’s self-confidence reward by re-noising the model’s own outputs and rewarding accurate recovery of the injected noise - a training-aligned signal that serves as a principled self-confidence score.
  • Across standard benchmarks and a comprehensive user study, ARC yields consistent gains in compositionality, text rendering, and text–image alignment, while modestly improving human-preference metrics.
  • ARC complements external-reward pipelines: applying ARC on top of externally post-trained models improves non-target capabilities (compositionality, text rendering, alignment) with only mild trade-offs on the targeted external metric, mitigating reward hacking.
@@ 3.1 Flow Matching and Rectified Flow
Figure 2: Overview of ARC. Given a text prompt $c$, we generate $G$ different latents. Without decoding, we re-noise the latents using $K$ noise probes across $t\in\mathcal{T}\subset[0,1]$. For each generated latent $z_{0}^{(i)}$, we formulate the text-to-image generative model’s self-confidence of the generated latent as the ability to denoise the re-noised latent. We leverage this self-confidence as an internal reward scalar value, which we use to post-train the text-to-image generative model using GRPO [59, 37]. We omit the KL term in this figure for better readability.
@@ 3.2 GRPO for Flow Matching
4 Adaptive Rewarding by Self-Confidence
Overview. We present ARC (Adaptive Rewarding by Self-Confidence), a post-training method for text-to-image generators that eschews reliance on external reward models. ARC leverages the model’s own self-confidence as an intrinsic reward: after generating an output, we re-noise it at selected timesteps and score how accurately the model recovers the injected noise. Aggregating these per-timestep recovery errors yields a single on-policy scalar reward for reinforcement learning. In the following, we detail the computation of the self-confidence reward (Sec. 4.1), and the stabilization and efficiency techniques used for using ARC (Sec. 4.2). An overview of ARC is shown in Fig. 2.
@@ 4.1 Intrinsic Self-Confidence Reward
$R_{\mathrm{ARC}}\big(z_{0}^{(i)},c\big)\;=\;\frac{1}{\sum_{t\in\mathcal{T}}w(t)}\sum_{t\in\mathcal{T}}w(t)\,S_{i,t}.$ (14)
We use $w(t)=1$ in practice for simplicity. Note that external rewards typically operate in pixel space, $R_{\mathrm{ext}}(x^{(i)},c)$, where $x^{(i)}=\mathrm{Dec}(z_{0}^{(i)})$ for a fixed decoder $\mathrm{Dec}\!:\mathcal{Z}\!\to\!\mathcal{X}$. In contrast, $R_{\mathrm{ARC}}$ is computed directly in latent space, avoiding decoding and keeping the signal model-native.
@@ 4.2 Stabilization and Efficiency Techniques
Self-Confidence calculation with CFG. Although $G$ images are sampled with CFG for GRPO training, our ARC self-confidence is computed without CFG. CFG forms a mixture field $v_{\mathrm{cfg}}=v_{\mathrm{uncond}}+s\,(v_{\mathrm{cond}}-v_{\mathrm{uncond}})$; this mixture would calculate a self-certainty based on the guided proxy rather than the base conditional model trained on data. In our experiments, omitting CFG during self-confidence computation yields stronger and more stable improvements.
Online calculation of self-confidence. When computing self-confidence for ARC, we can either compute it (1) online, using the model being trained ($\pi_{\theta}$), or (2) offline, using a fixed based model ($\pi_{\textrm{ref}}$) for the self-confidence calculation. While we do not observe severe over-optimization [20] when computing the self-confidence in an offline manner, we notice that the performance improves when we compute self-confidence rewards in an online manner. We conjecture that as the model improves through ARC post-training, the stability and reliability of self-confidence improves as well, leading to a stronger improvement in performance.
@@
Table 1: Quantitative results of ARC. We evaluate ARC post-training on base models of SD3.5 [16], on GenEval [21], Text Rendering, human preference reward models [30, 75, 78, 72], and image quality metrics. We show that ARC post-training yields consistent gains across text-to-image generative models on different quantitative metrics.
Figure 3: User study against baseline SD3.5-M [16] on PartiPrompts [56] and HPSv2 [75]. The user study shows that ARC post-training yields favorable visual realism/appeal, and text-image alignment.
@@ 5.3 Results
Quantitative results. The evaluation results are shown in Tab. 1, where it can be seen that applying our proposed ARC on the T2I flow-matching model SD3.5-M for post-training yields consistent gains across task-specific, image quality, and human preference metrics. While improvements in human preference is modest, we observe a considerable improvement in performance on compositional generation (GenEval [21]), text rendering (OCR[14]) and CLIPScore [53], almost matching the performance of SD3.5-L in these three metrics, albeit having less than $\frac{1}{3}$ of the parameters (2.5B vs. 7.1B). This evidences that the text-to-image generative model’s intrinsic confidence is strongly correlated with the success of compositional generation, text rendering and text-image alignment, which are more objective metrics compared to human preferences.
Figure 4: Effect of ARC post-training SD3.5-M after post-training on PickScore [30] using FlowGRPO [37]. ARC complements external rewards, showing the best best compositional generation and visual appeal on GenEval [21]. Post-training on external rewards yields high visual appeal, but sacrifices compositionality as shown above (Column 3: Generates yellow motorcycle instead / generates unwanted human).
Figure 5: Qualitative results of ARC when applied on SD3.5 [16] on DrawBench [56], GenEval [21] and OCR [14]. It can be seen that applying ARC shows consistent improvements over the baseline SD3.5.
@@
Table 2: Ablation study results of ARC. We validate the design choices of ARC over number of noise probes $K$, the usage of CFG for self-confidence calculation, and online/offline self-confidence calculation. Our current configurations yield superior results.
We also analyze the effect of applying ARC post-training after post-training SD3.5-M on external rewards using Flow-GRPO [37]. The results show that while the performance on the external reward itself is mildly compromised, we consistently gain noticeable improvements across GenEval, OCR and CLIPSCcore, which strengthens our hypothesis that the intrinsic confidence of text-to-image generative model is strongly correlated with with the success of compositional generation, text rendering and text-image alignment. In Fig. 4, we provide visual examples of SD3.5-M post-trained with FlowGRPO using PickScore as the external reward, finally post-trained with ARC intrinsic self-certainty rewards. It shows that integrating ARC improves the generative performances even further, evidencing the complementarity of ARC and external rewards.
User study. In Fig. 3, we provide the results of a user study on prompts from PartiPrompt [81] and HPSv2 [75], asking users to assess the generated images based on visual appeal/realism and text alignment. We summarize $\sim$1,800 responses from 20 participants. The user study results show that applying ARC post-training SD3.5-M consistently outperforms the baseline model in terms of visual realism/appeal and text alignment.
Qualitative comparison. We provide additional qualitative comparisons in Fig. 1 and Fig. 5, where it can be seen that applying ARC yields visually pleasurable results with improved compositionality and text rendering capabilities, even when trained with no external reward at all.
@@ 5.4 Ablation study and analyses
In Tab. 2, we provide ablation study results to validate the design and hyperparameter choices of ARC.
@@
5.5 Limitations
One limitation of ARC is that the intrinsic self-confidence does not align strongly with human preference, where observed gains are modest. Also, while ARC shows to improve compositional generation, text rendering and text faithfulness, we cannot target a specific alignment objective using ARC alone. However, we showed that ARC can be integrated with external rewards to target specific alignments while alleviating reward hacking and improving compositionality of text rendering capabilities (Tab. 1).
@@ 6 Conclusion
We introduced ARC, a reinforcement-learning post-training framework that replaces external critics with an intrinsic self-confidence signals. ARC converts this self-confidence into stable scalar rewards and optimizes a flow-matching text-to-image generator without auxiliary annotators or task-specific validators. Across benchmarks and a user study, ARC yields consistent improvements in compositionality, text rendering and text alignment; a comprehensive user study supports our findings. We also showed that ARC can be integrated with external rewards to align text-to-image generation to specific external criteria with mitigated reward hacking and additional improvements. Future directions include (i) consistency-aware extensions (temporal/multi-view) to carry ARC to video and 3D generation, and (ii) disentangling and calibrating intrinsic signals to enable precise, task-targeted reward shaping.
[…]
@@ 7 Rationale of ARC
We motivate ARC by testing whether a model’s denoising-based self-confidence correlates with perceived image quality when generation settings are made progressively stronger while keeping the underlying model fixed. Specifically, we compare three inference regimes for the same text–to–image model: (i) $10$ sampling steps without CFG, (ii) $10$ steps with CFG, and (iii) $20$ steps with CFG in Fig. 6. For each regime, we compute the self-confidence score defined in Sec. 4.1 on the model’s own outputs (same scorer and parameters across all cases). Empirically, the self-confidence distribution increases from (i) to (ii) to (iii), matching the observed rise in visual fidelity and prompt adherence. Because the signal is computed post hoc by the same model regardless of guidance or step count, these trends are not artifacts of an external evaluator; rather, they indicate that better samples are also easier for the model to self-denoise. This observation provides the rationale for treating denoising-based self-confidence as a model-native reward for post-training.
@@
Figure 6: Rationale of ARC. Distributions of the denoising-based self-confidence under three inference settings—$10$ steps (no CFG), $10$ steps (CFG), and 2020 steps (CFG). The distribution shifts monotonically rightward (higher self-confidence) in the same order that visual quality improves, indicating that the ability to recover injected noise is predictive of sample quality even when the scorer is the same model. This alignment underpins ARC’s use of self-confidence as an intrinsic reward.
8 ARC Post-Training on SD3.5-L
@@
To assess scalability, we apply ARC to SD3.5-L [16], a larger base model than the SD3.5-M used in the main experiments. Unless otherwise noted, we reuse the same training recipe (shortened denoising horizon, suffix-only updates, shared probes, CFG-free scoring). As reported in Tab. 3, ARC yields consistent gains in compositional generation, text rendering, and text–image alignment, while remaining competitive on human-preference metrics (e.g., HPSv2, PickScore). These results suggest that ARC scales to higher-capacity text-to-image models without inducing reward hacking and remains effective beyond the SD3.5-M setting.
9 Applying ARC on FLUX.1-Dev
@@
To test architectural generality, we apply ARC to FLUX.1 - Dev [3], a flow-matching text-to-image generator with a design distinct from SD3.5. We keep the core ARC recipe unchanged (shortened denoising horizon, suffix-only updates, shared probes, CFG-free scoring), adapting only to the model’s native scheduler and inference step count. A small deviation is the suffix window: we set $\rho=0.5$, i.e., train on the latter half of the scheduler steps, which increased training stability in this setting. As reported in Tab. 3, ARC delivers consistent gains in compositional generation, text rendering, and text–image alignment, while remaining competitive on human-preference metrics (e.g., HPSv2, PickScore). The results indicate that ARC transfers effectively across architectures and remains robust on another representative flow-matching T2I model.
@@
+ ARC (Ours) 0.66 0.65 0.271 5.67 22.69 0.292 0.90 3.23
Table 3: Applying ARC to SD3.5-L [16] and FLUX.1-Dev [3]. We apply ARC on additional models of SD3.5-L and FLUX.1-Dev, to verify the effect of ARC given (1) a larger base model, and (2) a different architecture from SD3.5-M. $\dagger$ denotes results taken from DiffusionNFT [91]. We base our experiments on our reproduced results based on the official weights of SD3.5-L [16] and FLUX.1-Dev [3]. The results show that ARC consistently results in improved compositionality, text rendering and text-image alignment, while being competitive at human preference metrics.

### v2

# Improving Text-to-Image Generation with Intrinsic Self-Confidence Rewards

@@ Abstract
Text-to-image generation powers content creation across design, media, and data augmentation. Post-training of text-to-image generative models is a promising path to better match human preferences, factuality, and improved aesthetics. We introduce SOLACE (Self-Originating LAtent Confidence Estimation), a post-training framework that replaces external reward supervision with an internal self-confidence signal, obtained by evaluating how accurately the model recovers injected noise under self-denoising probes. SOLACE converts this intrinsic signal into scalar rewards, enabling fully unsupervised optimization without additional datasets, annotators, or reward models. Empirically, by reinforcing high-confidence generations, SOLACE delivers consistent gains in compositional generation, text rendering and text-image alignment over the baseline. We also find that integrating SOLACE with external rewards results in a complementary improvement, with alleviated reward hacking.
Figure 1: Qualitative examples of SOLACE on Pick-a-Pic dataset [30]. Best viewed on electronics.
@@ 1 Introduction
Despite extensive progress in extrinsically supervised post-training, relying on intrinsic signals remains under-explored for text-to-image generation. In this work, we aim to answer a fundamental question: can internal feedback from the text-to-image generator itself provide meaningful signals for post-training? To this end, we introduce Self-Originating LAtent Confidence Estimation (SOLACE), a post-training framework that relies on intrinsic self-certainty as the model-native reward. Inspired by Score Distillation Sampling [50, 62, 74], which leverages a pretrained text-to-image generator as a critic for text-to-3D or -4D generation, we propose to leverage a text-to-image generator to critique its own generation. Concretely, given a sampled latent $z_{0}$, we re-noise $z_{0}$ to selected timesteps $t\in\mathcal{T}$ using the model’s forward noising schedule, and measure how well the model reconstructs the injected noise, rewarding small residuals as ’high confidence’. Our hypothesis is that large-scale pretraining in diffusion generative models endows them with priors over real images and text–image alignment, so the model’s self-confidence would align strongly with text-alignment and realism.
Empircally, SOLACE yields consistent gains in compositional generation [21], text rendering [14], and text–image alignment [53], while modestly improving human-preference scores [30, 75, 78, 72] without external rewards. Qualitative side-by-side comparisons and a comprehensive user study corroborate these trends. Together, the results indicate that intrinsic self-confidence aligns with key facets of image generation. Moreover, applying SOLACE on top of an extrinsically post-trained model (i.e., using external rewards) recovers additional improvements in compositionality, text rendering, and alignment, with only slight drops on the targeted external metric, evidencing that intrinsic and extrinsic rewards are complementarily beneficial for post-training stronger T2I generators.
@@
  • We present SOLACE (Self-Originating LAtent Confidence Estimation), a post-training framework to use self-confidence as a reward for text-to-image generators.
  • We define SOLACE’s self-confidence reward by re-noising the model’s own outputs and rewarding accurate recovery of the injected noise - a training-aligned signal that serves as a principled self-confidence score.
  • Across standard benchmarks and a comprehensive user study, SOLACE yields consistent gains in compositionality, text rendering, and text–image alignment, while modestly improving human-preference metrics.
  • SOLACE complements external-reward pipelines: applying SOLACE on top of externally post-trained models improves non-target capabilities (compositionality, text rendering, alignment) with only mild trade-offs on the targeted external metric, mitigating reward hacking.
@@ 3.1 Flow Matching and Rectified Flow
Figure 2: Overview of SOLACE. Given a text prompt $c$, we generate $G$ different latents. Without decoding, we re-noise the latents using $K$ noise probes across $t\in\mathcal{T}\subset[0,1]$. For each generated latent $z_{0}^{(i)}$, we formulate the text-to-image generative model’s self-confidence of the generated latent as the ability to denoise the re-noised latent. We leverage this self-confidence as an internal reward scalar value, which we use to post-train the text-to-image generative model using GRPO [59, 37]. We omit the KL term in this figure for better readability.
@@ 3.2 GRPO for Flow Matching
4 Self-Originating LAtent Confidence Estimation
Overview. We present SOLACE (Self-Originating LAtent Confidence Estimation), a post-training method for text-to-image generators that eschews reliance on external reward models. SOLACE leverages the model’s own self-confidence as an intrinsic reward: after generating an output, we re-noise it at selected timesteps and score how accurately the model recovers the injected noise. Aggregating these per-timestep recovery errors yields a single on-policy scalar reward for reinforcement learning. In the following, we detail the computation of the self-confidence reward (Sec. 4.1), and the stabilization and efficiency techniques used for using SOLACE (Sec. 4.2). An overview of SOLACE is shown in Fig. 2.
@@ 4.1 Intrinsic Self-Confidence Reward
$R_{\mathrm{SOLACE}}\big(z_{0}^{(i)},c\big)\;=\;\frac{1}{\sum_{t\in\mathcal{T}}w(t)}\sum_{t\in\mathcal{T}}w(t)\,S_{i,t}.$ (14)
We use $w(t)=1$ in practice for simplicity. Note that external rewards typically operate in pixel space, $R_{\mathrm{ext}}(x^{(i)},c)$, where $x^{(i)}=\mathrm{Dec}(z_{0}^{(i)})$ for a fixed decoder $\mathrm{Dec}\!:\mathcal{Z}\!\to\!\mathcal{X}$. In contrast, $R_{\mathrm{SOLACE}}$ is computed directly in latent space, avoiding decoding and keeping the signal model-native.
@@ 4.2 Stabilization and Efficiency Techniques
Self-Confidence calculation with CFG. Although $G$ images are sampled with CFG for GRPO training, our SOLACE self-confidence is computed without CFG. CFG forms a mixture field $v_{\mathrm{cfg}}=v_{\mathrm{uncond}}+s\,(v_{\mathrm{cond}}-v_{\mathrm{uncond}})$; this mixture would calculate a self-certainty based on the guided proxy rather than the base conditional model trained on data. In our experiments, omitting CFG during self-confidence computation yields stronger and more stable improvements.
Online calculation of self-confidence. When computing self-confidence for SOLACE, we can either compute it (1) online, using the model being trained ($\pi_{\theta}$), or (2) offline, using a fixed based model ($\pi_{\textrm{ref}}$) for the self-confidence calculation. While we do not observe severe over-optimization [20] when computing the self-confidence in an offline manner, we notice that the performance improves when we compute self-confidence rewards in an online manner. We conjecture that as the model improves through SOLACE post-training, the stability and reliability of self-confidence improves as well, leading to a stronger improvement in performance.
@@
Table 1: Quantitative results of SOLACE. We evaluate SOLACE post-training on base models of SD3.5 [16], on GenEval [21], Text Rendering, human preference reward models [30, 75, 78, 72], and image quality metrics. We show that SOLACE post-training yields consistent gains across text-to-image generative models on different quantitative metrics.
Figure 3: User study against baseline SD3.5-M [16] on PartiPrompts [56] and HPSv2 [75]. The user study shows that SOLACE post-training yields favorable visual realism/appeal, and text-image alignment.
@@ 5.3 Results
Quantitative results. The evaluation results are shown in Tab. 1, where it can be seen that applying our proposed SOLACE on the T2I flow-matching model SD3.5-M for post-training yields consistent gains across task-specific, image quality, and human preference metrics. While improvements in human preference is modest, we observe a considerable improvement in performance on compositional generation (GenEval [21]), text rendering (OCR[14]) and CLIPScore [53], almost matching the performance of SD3.5-L in these three metrics, albeit having less than $\frac{1}{3}$ of the parameters (2.5B vs. 7.1B). This evidences that the text-to-image generative model’s intrinsic confidence is strongly correlated with the success of compositional generation, text rendering and text-image alignment, which are more objective metrics compared to human preferences.
Figure 4: Effect of SOLACE post-training SD3.5-M after post-training on PickScore [30] using FlowGRPO [37]. SOLACE complements external rewards, showing the best best compositional generation and visual appeal on GenEval [21]. Post-training on external rewards yields high visual appeal, but sacrifices compositionality as shown above (Column 3: Generates yellow motorcycle instead / generates unwanted human).
Figure 5: Qualitative results of SOLACE when applied on SD3.5 [16] on DrawBench [56], GenEval [21] and OCR [14]. It can be seen that applying SOLACE shows consistent improvements over the baseline SD3.5.
@@
Table 2: Ablation study results of SOLACE. We validate the design choices of SOLACE over number of noise probes $K$, the usage of CFG for self-confidence calculation, and online/offline self-confidence calculation. Our current configurations yield superior results.
We also analyze the effect of applying SOLACE post-training after post-training SD3.5-M on external rewards using Flow-GRPO [37]. The results show that while the performance on the external reward itself is mildly compromised, we consistently gain noticeable improvements across GenEval, OCR and CLIPSCcore, which strengthens our hypothesis that the intrinsic confidence of text-to-image generative model is strongly correlated with with the success of compositional generation, text rendering and text-image alignment. In Fig. 4, we provide visual examples of SD3.5-M post-trained with FlowGRPO using PickScore as the external reward, finally post-trained with SOLACE intrinsic self-certainty rewards. It shows that integrating SOLACE improves the generative performances even further, evidencing the complementarity of SOLACE and external rewards.
User study. In Fig. 3, we provide the results of a user study on prompts from PartiPrompt [81] and HPSv2 [75], asking users to assess the generated images based on visual appeal/realism and text alignment. We summarize $\sim$1,800 responses from 20 participants. The user study results show that applying SOLACE post-training SD3.5-M consistently outperforms the baseline model in terms of visual realism/appeal and text alignment.
Qualitative comparison. We provide additional qualitative comparisons in Fig. 1 and Fig. 5, where it can be seen that applying SOLACE yields visually pleasurable results with improved compositionality and text rendering capabilities, even when trained with no external reward at all.
@@ 5.4 Ablation study and analyses
In Tab. 2, we provide ablation study results to validate the design and hyperparameter choices of SOLACE.
@@
5.5 Limitations of SOLACE
One limitation is that the intrinsic self-confidence does not align strongly with human preference, where observed gains are modest. Also, while SOLACE shows to improve compositional generation, text rendering and text faithfulness, we cannot target a specific alignment objective using SOLACE alone. However, we showed that SOLACE can be integrated with external rewards to target specific alignments while alleviating reward hacking and improving compositionality or text rendering capabilities (Tab. 1).
@@ 6 Conclusion
We introduced SOLACE, a reinforcement-learning post-training framework that replaces external critics with an intrinsic self-confidence signals. SOLACE converts this self-confidence into stable scalar rewards and optimizes a flow-matching text-to-image generator without auxiliary annotators or task-specific validators. Across benchmarks and a user study, SOLACE yields consistent improvements in compositionality, text rendering and text alignment; a comprehensive user study supports our findings. We also showed that SOLACE can be integrated with external rewards to align text-to-image generation to specific external criteria with mitigated reward hacking and additional improvements. Future directions include (i) consistency-aware extensions (temporal/multi-view) to carry SOLACE to video and 3D generation, and (ii) disentangling and calibrating intrinsic signals to enable precise, task-targeted reward shaping.
Acknowledgement. This work was supported by the IITP grants (RS-2022-II220290: Visual Intelligence for Space-Time Understanding and Generation based on Multi-layered Visual Common Sense (40%), RS-2022-II220113: Developing a Sustainable Collaborative Multi-modal Lifelong Learning Framework (50%), RS-2019-II191906: AI Graduate School Program at POSTECH (5%), RS-2025-02653113: High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale (5%)) funded by the Korea government (MSIT).
@@ 7 Rationale of SOLACE
We motivate SOLACE by testing whether a model’s denoising-based self-confidence correlates with perceived image quality when generation settings are made progressively stronger while keeping the underlying model fixed. Specifically, we compare three inference regimes for the same text–to–image model: (i) $10$ sampling steps without CFG, (ii) $10$ steps with CFG, and (iii) $20$ steps with CFG in Fig. 6. For each regime, we compute the self-confidence score defined in Sec. 4.1 on the model’s own outputs (same scorer and parameters across all cases). Empirically, the self-confidence distribution increases from (i) to (ii) to (iii), matching the observed rise in visual fidelity and prompt adherence. Because the signal is computed post hoc by the same model regardless of guidance or step count, these trends are not artifacts of an external evaluator; rather, they indicate that better samples are also easier for the model to self-denoise. This observation provides the rationale for treating denoising-based self-confidence as a model-native reward for post-training.
@@
Figure 6: Rationale of SOLACE. Distributions of the denoising-based self-confidence under three inference settings—$10$ steps (no CFG), $10$ steps (CFG), and 2020 steps (CFG). The distribution shifts monotonically rightward (higher self-confidence) in the same order that visual quality improves, indicating that the ability to recover injected noise is predictive of sample quality even when the scorer is the same model. This alignment underpins SOLACE’s use of self-confidence as an intrinsic reward.
8 SOLACE Post-Training on SD3.5-L
@@
To assess scalability, we apply SOLACE to SD3.5-L [16], a larger base model than the SD3.5-M used in the main experiments. Unless otherwise noted, we reuse the same training recipe (shortened denoising horizon, suffix-only updates, shared probes, CFG-free scoring). As reported in Tab. 3, SOLACE yields consistent gains in compositional generation, text rendering, and text–image alignment, while remaining competitive on human-preference metrics (e.g., HPSv2, PickScore). These results suggest that SOLACE scales to higher-capacity text-to-image models without inducing reward hacking and remains effective beyond the SD3.5-M setting.
9 Applying SOLACE on FLUX.1-Dev
@@
To test architectural generality, we apply SOLACE to FLUX.1 - Dev [3], a flow-matching text-to-image generator with a design distinct from SD3.5. We keep the core SOLACE recipe unchanged (shortened denoising horizon, suffix-only updates, shared probes, CFG-free scoring), adapting only to the model’s native scheduler and inference step count. A small deviation is the suffix window: we set $\rho=0.5$, i.e., train on the latter half of the scheduler steps, which increased training stability in this setting. As reported in Tab. 3, SOLACE delivers consistent gains in compositional generation, text rendering, and text–image alignment, while remaining competitive on human-preference metrics (e.g., HPSv2, PickScore). The results indicate that SOLACE transfers effectively across architectures and remains robust on another representative flow-matching T2I model.
@@
+ SOLACE (Ours) 0.66 0.65 0.271 5.67 22.69 0.292 0.90 3.23
Table 3: Applying SOLACE to SD3.5-L [16] and FLUX.1-Dev [3]. We apply SOLACE on additional models of SD3.5-L and FLUX.1-Dev, to verify the effect of SOLACE given (1) a larger base model, and (2) a different architecture from SD3.5-M. $\dagger$ denotes results taken from DiffusionNFT [91]. We base our experiments on our reproduced results based on the official weights of SD3.5-L [16] and FLUX.1-Dev [3]. The results show that SOLACE consistently results in improved compositionality, text rendering and text-image alignment, while being competitive at human preference metrics.